OpenStudy (anonymous):

An isosceles triangle has area of 140 ft2. If the base is 10 ft, what is the length of each leg? Round your answer to the nearest tenth.

OpenStudy (anonymous):

What is the formula for the area of a triangle? And what does it mean when a triangle is isosceles?

OpenStudy (anonymous):

A= 1/2bh and it means it has 2 equal sides

OpenStudy (anonymous):

In the question, A = 140. Therefore, the total area of the triangle is 140: (1/2)bh = 140. The base is given as 10: (1/2)(10)h = 140 The value for h can be found by solving the equation. Then, you have a right angled triangle and you can use Pythagoras to find the length of the longest side.

OpenStudy (anonymous):

The longest side of the right angled triangle will be the same length of one of the equal sides of the original triangle.